find a 3-by-3 magic square using the numbers 6,7,8,14,15,16,22,23 and 24


7 _ _
_ _ _
_ 8 _

Answer by Edwin McCravy(20056)

Let the unknowns be A,B,C,D,E,F,G, like this 

7 A B
C D E
F 8 G

I will number my equations in parentheses:

We sum all the numbers that go in the magic square:

(1) 6+7+8+14+15+16+22+23+24=135

therefore:

Each row, column, and diagonal must sum to 1/3 of 135 or 45

(2) 7+A+B = C+D+E = F+8+G = 7+C+F = A+D+8 = B+E+G = 7+D+G = B+D+F = 45

From (2) we get a system of 8 equations in 7 unknowns:

(3)  A + B                      = 38
(4)          C + D + E          = 45
(5)                       F + G = 37
(6)          C          + F     = 38
(7)  A         + D              = 37
(8)     B          + E      + G = 45
(9)              D          + G = 38
(10)    B      + D      + F     = 45      

Eliminate A by subtracting (3)-(7), getting (11)


(4)          C + D + E          = 45
(5)                       F + G = 37
(6)          C          + F     = 38
(8)     B          + E      + G = 45
(9)              D          + G = 38
(10)    B      + D      + F     = 45
(11)    B      - D              =  1

Eliminate C by subtracting (4)-(6) getting (12)

(5)                       F + G = 37
(8)     B          + E      + G = 45
(9)              D          + G = 38
(10)    B      + D      + F     = 45
(11)    B      - D              =  1
(12)             D + E  - F     =  7

Eliminate E by subtracting (8)-(12) getting (13)

(5)                       F + G = 37
(9)              D          + G = 38
(10)    B      + D      + F     = 45
(11)    B      - D              =  1
(13)    B      - D      + F + G = 38

Eliminate D by adding (9)+(10)+(11)+(13) getting (14)

(5)                      F +  G  =  37
(14)   3B             + 2F + 2G  = 122

We eliminate F and G by multiplying (5) by -2 getting (15)

(15)                 - 2F - 2G  =  -74 

and adding to (14) 

(15)                  - 2F - 2G  = -74
(14)   3B             + 2F + 2G  = 122
--------------------------------------
       3B                        =  48 

so B = 16 

Substituting in (11)           Substituting in (3)
(11)    B - D = 1                      A + B = 38
       16 - D = 1                     A + 16 = 38 
           -D = -15                        A = 22
            D = 15

Substituting in (9)
(9)      D + G = 38
        15 + G = 38
             G = 23

Substituting in (5)
(5)     F + G = 37
       F + 23 = 37
            F = 14

Substituting in (6)
(6)     C + F = 38
       C + 14 = 38
            C = 24

Substituting C = 24 and D = 15 in (4)
(4)  C + D + E  = 45
   24 + 15 + E = 45 
        39 + E = 45
             E =  6 

So the magic square

7 A B     7 22 16
C D E  = 24 15  6 
F 8 G    14  8 23

Edwin